OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 9, 15..., if there are 36 terms?

OpenStudy (welshfella):

Sum of n terms = (n/2)[2a + d(n-1)] where d = common difference and a = first term

OpenStudy (welshfella):

common difference d = 9- 3 = 6

OpenStudy (anonymous):

right then what do I do?

OpenStudy (welshfella):

well plug in the values n = 36, d = 6 and a = 3

OpenStudy (welshfella):

plug these in and work it out

OpenStudy (welshfella):

= (36/2)[2*3 - 6(36-1)]

OpenStudy (welshfella):

sorry that should be a + before the 6 not -

OpenStudy (anonymous):

you're all good so then I have to solve that?

OpenStudy (welshfella):

do the parts in the parentheses first

OpenStudy (anonymous):

(18)[35] is what I got but that doesn't seem like it makes sense at all

OpenStudy (welshfella):

36/2 = 18 36 - 1 = 35

OpenStudy (welshfella):

no thats not right its 18(6 + 6*35) do the stuff in the parentheses first then finall y multiply by 18

OpenStudy (anonymous):

7560?

OpenStudy (welshfella):

multiply 6 by 35 before adding the 6

OpenStudy (anonymous):

actually 3888

OpenStudy (welshfella):

no i think you are doing things in the wrong order remember PEDMAS Parentheses Exponent Division Multiplication Add Subtract thats the order of operations

OpenStudy (welshfella):

yes 3888 is correct 18(6 + 6*35 = 18*(6 + 210) = 18 * 216 = 3888

OpenStudy (anonymous):

thank you!

OpenStudy (welshfella):

PEDMAS is very important

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

note i worked out 6 + 210 before i multipilied by 18 because 6 + 210 was in the parentheses